# DNABERT-S: Learning Species-Aware DNA Embedding with Genome Foundation Models

### v1

@@ C.3 Impact of Sequence Length
C.3.1 Varying Sequence Length in Training
Training with longer DNA sequences increases the need for more memory and computing power. It also means we can only use smaller batches of data at a time. Therefore, the length of the sequences is an important factor in contrastive training as it affects how much it costs to train the model. To see how different sequence lengths affect training, we did three experiments using the same data. Our training data has sequences that are 1000010000bp long. For experiments with shorter sequences $S$, we only used the first $S$ nucleotides of each DNA sequence. We tested sequence lengths of 500500bp, 20002000bp, and 1000010000bp, training only with Weighted SimCLR loss and starting from the pre-trained DNABERT-2 model.
Figure 6 shows the results for the three models, along with the pre-trained DNABERT-2 without contrastive training and the strongest baseline, TNF. The findings reveal that sequence length significantly influences the model’s performance. Training even on short sequences, such as 500500bp, leads to substantial improvements. The model trained with 500500bp sequences performs nearly as well as TNF. When we increase the input sequence length from 500500bp to 20002000bp, there’s a marked improvement in performance. A similar trend is observed when increasing the sequence length from 20002000bp to 1000010000bp. These results highlight the importance of sequence length in training an effective model. Therefore, we decided to train our model with 1000010000bp sequences, despite the higher computational requirements.
@@ C.3.1 Varying Sequence Length in Training
C.3.2 Varying Sequence Length in Evaluation
In this part, we assess how the length of DNA sequences in evaluation impacts performance. We use two synthetic datasets for clustering and classification tasks. Each sequence in these datasets is deliberately constructed to be 1000010000bp long. This allows us to create a test set where all sequences have the same length. We test sequence lengths ranging from 3232bp ($2^{5}$) to 81928192bp ($2^{13}$). For each test with different sequence lengths, we keep everything else the same, like how we split the data into training and testing sets and the settings for logistic regression.
Figure 7 presents the performance of TNF and DNABERT-S with various sequence lengths. The results show that both models significantly benefit from longer sequences. When the sequence length is less than 256256bp ($2^{8}$), both models perform poorly in clustering and classifying samples. However, as the sequence length increases, starting from 512512bp ($2^{9}$), DNABERT-S begins to outperform TNF. The performance gap between the two models gets bigger as the sequence length increases. These findings highlight the crucial role of sequence length in effectively differentiating between species.
C.4 Impact of Embedding Dimension Reduction
@@
Figure 8 illustrates DNABERT-S’s performance with these varying feature dimensions, in comparison to TNF. The results demonstrate that DNABERT-S’s embedding is quite resilient to dimension compression. It maintains nearly the same performance level even when reduced to 256 dimensions and only experiences a notable drop in performance when compressed to 32 dimensions. Remarkably, DNABERT-S still surpasses the 256-dimensional TNF feature even when its own dimensionality is reduced to just 16. This robustness to dimension reduction enhances its practical applicability in various genomic contexts.
@@
C.5 Results on Scenario with Abundant Labeled Data
While previous experiments have highlighted DNABERT-S’s exceptional performance in scenarios with limited or no labeled data, this section focuses on its effectiveness in situations where abundant labeled data is available.
We utilize the fungi and viral species classification datasets from Zhou et al. [2023]. The fungi dataset contains 20 different species, each with 400 training samples that are 10,000 base pairs (1000010000bp) long. The viral dataset comprises 40 different species, with each species providing 100 training samples of 5,000 base pairs (50005000bp) in length. Similar to our approach in the few-shot classification evaluation, we use DNABERT-2 and DNABERT-S to generate embeddings for each DNA sequence and subsequently train a logistic regression model for classification purposes. In alignment with Zhou et al. [2023], we employ the Matthews Correlation Coefficient (MCC) as our evaluation metric.
Table 6 presents the performance comparison between DNABERT-2 and DNABERT-S. As indicated by the results, DNABERT-S continues to demonstrate significant performance improvements even in the presence of a large volume of training data. This suggests that DNABERT-S’s capabilities are not confined to scenarios with scarce data but also extend effectively to data-rich environments, reinforcing its versatility and robustness in various genomic analysis contexts.
Model Fungi Viral
DNABERT-2 $91.37$ $35.44$
DNABERT-S 94.64 41.91
Table 6: Experimental results on the fungi and species classification dataset on GUE [Zhou et al., 2023]. DNABERT-S still outperforms DNABERT-2 when labeled data is abundant.
C.6 Results on Other Types of Tasks
This section evaluates DNABERT-S’s performance across a range of distinct genomic tasks. We utilize the GUE benchmark [Zhou et al., 2023], which comprises a comprehensive collection of 28 datasets covering 7 diverse tasks, such as epigenetic marks prediction, promoter prediction, and transcription factor binding site prediction. Following our established methodology, DNABERT-2 and DNABERT-S are used to generate embeddings for each DNA sequence, and a logistic regression model is trained for classification. The Matthews Correlation Coefficient (MCC) serves as the evaluation metric.
@@ C.6 Results on Other Types of Tasks
Table 7: Performance of DNABERT-2 and DNABERT-S on the GUE benchmark.
The results, as detailed in Table 7, show that DNABERT-2 outperforms DNABERT-S on 19 of the 28 datasets, with an average improvement of 2.12 in the MCC. This outcome aligns with expectations, given the distinct objectives of DNABERT-S and the GUE benchmark. DNABERT-S is specifically designed to generate species-aware embeddings, which are optimized for clustering and differentiating various species. In contrast, the tasks in the GUE benchmark, which predominantly involve predicting functional aspects of human and mouse genomes, require a different set of embedding characteristics. Therefore, it’s evident that DNABERT-S, while not universally superior in all aspects of DNA embedding generation, excels in its targeted domain. The training approach and methodologies developed for DNABERT-S, despite the model’s specialized focus, hold potential applicability across various genomic analysis scenarios

### v3

@@ C.6 Impact of Sequence Length
C.6.1 Varying Sequence Length in Training
Training with longer DNA sequences increases the need for more memory and computing power. It also means we can only use smaller batches of data at a time. Therefore, the length of the sequences is an important factor in contrastive training as it affects how much it costs to train the model. To see how different sequence lengths affect training, we did three experiments using the same data. Our training data has sequences that are $10000$bp long. For experiments with shorter sequences $S$, we only used the first $S$ nucleotides of each DNA sequence. We tested sequence lengths of $500$bp, $2000$bp, and $10000$bp, training only with Weighted SimCLR loss and starting from the pre-trained DNABERT-2 model.
Figure 7 shows the results for the three models, along with the pre-trained DNABERT-2 without contrastive training and the strongest baseline, TNF. The findings reveal that sequence length significantly influences the model’s performance. Training even on short sequences, such as $500$bp, leads to substantial improvements. The model trained with $500$bp sequences performs nearly as well as TNF. When we increase the input sequence length from $500$bp to $2000$bp, there’s a marked improvement in performance. A similar trend is observed when increasing the sequence length from $2000$bp to $10000$bp. These results highlight the importance of sequence length in training an effective model. Therefore, we decided to train our model with $10000$bp sequences, despite the higher computational requirements.
Figure 7: Performance of DNABERT-S in clustering and classification with different sequence lengths used in contrastive training.
C.6.2 Varying Sequence Length in Evaluation
In this part, we assess how the length of DNA sequences in evaluation impacts performance. We use two synthetic datasets for clustering and classification tasks. Each sequence in these datasets is deliberately constructed to be $10000$bp long. This allows us to create a test set where all sequences have the same length. We test sequence lengths ranging from $32$bp ($2^{5}$) to $8192$bp ($2^{13}$). For each test with different sequence lengths, we keep everything else the same, like how we split the data into training and testing sets and the settings for logistic regression.
Figure 8 presents the performance of TNF and DNABERT-S with various sequence lengths. The results show that both models significantly benefit from longer sequences. When the sequence length is less than $256$bp ($2^{8}$), both models perform poorly in clustering and classifying samples. However, as the sequence length increases, starting from $512$bp ($2^{9}$), DNABERT-S begins to outperform TNF. The performance gap between the two models gets bigger as the sequence length increases. These findings highlight the crucial role of sequence length in effectively differentiating between species.
@@ C.6.2 Varying Sequence Length in Evaluation
C.7 Impact of Embedding Dimension Reduction
@@
Figure 9 illustrates DNABERT-S’s performance with these varying feature dimensions, in comparison to TNF. The results demonstrate that DNABERT-S’s embedding is quite resilient to dimension compression. It maintains nearly the same performance level even when reduced to 256 dimensions and only experiences a notable drop in performance when compressed to 32 dimensions. Notably, DNABERT-S still surpasses the 256-dimensional TNF feature even when its own dimensionality is reduced to just 16. This robustness to dimension reduction enhances its practical applicability in various genomic contexts.
@@
C.8 Species-Aware Embeddings on Genomics Function Predictions
This section evaluates the impact of species-aware embedding on other types of genomics analysis tasks. Specifically, we aim to understand how the species-aware embeddings perform compared to the embedding generated by the genome foundation model without contrastive training on various types of genome function prediction tasks. We utilize the GUE benchmark [Zhou et al., 2023], which comprises a comprehensive collection of 28 datasets covering 7 diverse tasks, such as epigenetic marks prediction, promoter prediction, and transcription factor binding site prediction. Following our established methodology, each model is respectively used to generate embeddings for each DNA sequence, and a logistic regression model is trained for classification. The Matthews Correlation Coefficient (MCC) serves as the evaluation metric. We perform experiments on both DNABERT-2 and HyenaDNA.
@@ C.8 Species-Aware Embeddings on Genomics Function Predictions
HyenaDNA w/o 23.48 58.28 54.88 21.34 22.15 30.50 28.71 46.59
HyenaDNA w/ 21.24 53.88 55.54 23.86 18.66 31.08 22.11 45.67
Table 12: Performance of DNABERT-2 and HyenaDNA before and after species-aware contrastive training on the GUE benchmark.
The results, as detailed in Table 12, show that after species-aware conservative training, DNABERT-S underperforms DNABERT-2 on 19 of the 28 datasets, with an average loss of 2.12 in the MCC. Similarly, on HyenaDNA, the one went through species-aware training underperforms the original one on 18 out of 28 datasets. As a model honed for species-aware tasks, DNABERT-S may exhibit reduced generalizability compared to broader genome foundation models (e.g., DNABERT-2). However, this specialized focus should not be viewed as a disadvantage. The intrinsic design of DNABERT-S—to prioritize species-specific features—may naturally limit its applicability to a broader range of tasks, a trade-off inherent to its specialized nature. For instance, when considering sequences from different species with varying functions, DNABERT-S’s training objective emphasizes species-specific similarities over functional commonalities, a choice that is deliberate for its targeted application.
Appendix D i-Mix for Original Contrastive Learning
The Weighted SimCLR method treat each instance $\{x_{i}\}_{i=1}^{B}$ and $\{x_{i^{+}}\}_{i=1}^{B}$ as anchors, with each anchor associated with $22B-$ 2 negative samples. This section explains why integrating the i-Mix method into Weighted SimCLR (Sec. 3.1) doubles memory or training time.
Moreover, for each anchor $x_{i}$, if i-Mix considers all samples from $\{x_{j}\}_{j=1}^{2B}\backslash\{x_{i}\}$ as either positive or negative samples, the model $f(\cdot)$ must process both the initial data instances $\{x_{i}\}_{i=1}^{2B}$ and mixed data instances $\{h_{i}^{0}\}_{i=1}^{2B}$ to generate their embeddings. Therefore, the i-Mix requires nearly twice more memory or training time compared to the method in Sec. 3.1 if using the same batch size.